OpenStudy (anonymous):

Write the logarithm in terms of the log of x, y, and z. log(cuberoot x^2)y(sqrt z^5)

OpenStudy (anonymous):

log(cuberoot x^2)y(sqrt z^5) =log[x^(2/3)y(z^(5/2))] =logx^(2/3)+logy+log(z^(5/2)) =(2/3)logx+logy+(5/2)logz

OpenStudy (anonymous):

my multiple choices are: (2/3)logx + (1/3)logz + (1/6)logy or 2logx + logy^1/2 + logz or (2/3)logx + (1/3)logy + (5/6)logz or or logx^2 + logy + (1/2)logz any suggestions since we can't match it?

OpenStudy (anonymous):

it didnt come out clear and may be cuberoot over the entire equation... if so I believe it may be the 3rd option